OpenStudy (anonymous):

determine wether the graphs of y=3x+6 and 2y+6x=4 are perpendicular

OpenStudy (anonymous):

no they are not. in the form y=mx+b, if m1*m2=-1, then the graphs are perpendicular. here, however, the second graph is y=-3x+2 -3*3=-9, not -1